OpenStudy (matthew120):

help with 19 questions

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What do you need help with

OpenStudy (matthew120):

i dont want help i want the direct answer plz its due by 3

OpenStudy (matthew120):

and iim not being rude srry

OpenStudy (anonymous):

Well you need to understand how to solve the questions if you can't solve any questions out of 19. Other than that, what are the questions?

OpenStudy (matthew120):

Which equation could be used to solve the problem? The Petersons' new car can go 32 miles on a gallon of gasoline. At that rate, how far could it go on 8 gallons (d)? A. d = 32 • 8 B. d = 32 ÷ 8 C. d = 8 ÷ 32 D. d = 32 + 8

OpenStudy (matthew120):

@rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

32/8 because 8 mile per gal

rebeccaxhawaii (rebeccaxhawaii):

Please click "best response" for the person that helped you the most, this just shows the appreciation for the time they took to assist. If you want further help with anymore questions, please fan me so in the long run you can see when I'm on. I'll be sure to take a look at your question to help :)

OpenStudy (matthew120):

so b

OpenStudy (anonymous):

What do you think Matthew120?